The sum of two numbers is 92. The difference of the two numbers is 20. What are the numbers?
Answer by mananth(16949)

x+y=92
x-y=20
add the two equations
2x=112
/2
x=56
x+y=92
plug
y+56=92
y=92-56
y=36